OpenStudy (anonymous):

Which is the equation of the line passing through the points A(3, -3) and B(-4, 11)?

OpenStudy (anonymous):

@TheAsker2002 @freckles @Whitemonsterbunny17 @tallymark124 @rishavraj @dan815

OpenStudy (anonymous):

@AlicaIsCray

OpenStudy (anonymous):

@Rachell_jpg

OpenStudy (anonymous):

Fan and medals at the ready

OpenStudy (anonymous):

is there a picture?

OpenStudy (anonymous):

Nope (word problems URG!!!)

OpenStudy (anonymous):

okay i still can help

OpenStudy (anonymous):

@Mateaus your awesome although any ideas math related?

OpenStudy (anonymous):

@Question_Asker @SamsungFanBoy @SavineTwist @BellaWolfe

OpenStudy (anonymous):

First you need to solve for the slope of the problem.

OpenStudy (anonymous):

okay, what is the best way to do that?

OpenStudy (anonymous):

formula to find slope where x and y values are (x1,y1)(x2,y2)

OpenStudy (anonymous):

okay so the slope is negative 2...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what would i need to do next

OpenStudy (anonymous):

Then you would plug in one of the original points as y and x and solve for b. Then rewrite the equation with the slope and the value you found for b.

OpenStudy (anonymous):

? wait what, im sorry what do i do?

OpenStudy (anonymous):

It's okay. Do you know what slope-intercept formula is? y=mx+b

OpenStudy (anonymous):

now i do lol

OpenStudy (anonymous):

Haha it's all good. Do you understand what to do?

OpenStudy (anonymous):

so plug in -2 as Y into (Y=mx+B )

OpenStudy (anonymous):

No. -2 is the slope. In y=mx+b, b is the y intercept and m is the slope. You would plug in -2 as m and then use one of the original points as x and y. Does that make sense?

OpenStudy (anonymous):

so it would look like Y=-2x=b?

OpenStudy (anonymous):

Not quite. It's supposed to be +b at the end. But other than that, yes.

OpenStudy (anonymous):

oh duh sorry

OpenStudy (anonymous):

It's okay. Then you plug in one of the two original points for x and y in order to solve for b.

OpenStudy (anonymous):

y = -2x + 3?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

that would be my final answer right

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thank you so much jessliynP

OpenStudy (anonymous):

that was a big one

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

Believe it our not im decent at math so message me any time i will do my best to return the favour :)

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Actually, do you happen to know how to find real or complex zeros of problems??

OpenStudy (anonymous):

are you talking about inside of a polynolial?

OpenStudy (anonymous):

ummmm.... maybe?? I think so.....

OpenStudy (anonymous):

the problem is f(x)=x^3+7x^2+9x-5 and I know that one of the zeros is x=-5

OpenStudy (anonymous):

https://www.youtube.com/watch?v=mSBT3oqOoJs (I watch this guy from time to time) Very helpful

OpenStudy (anonymous):

okay. thanks! i'll try this....

OpenStudy (anonymous):

im short on time so i wont explain but your answer can be.... 1) x = 5 2) x =(2-√8)/2=1-√ 2 = -0.414 3) x =(2+√8)/2=1+√ 2 = 2.414

OpenStudy (anonymous):

@JessilynP

OpenStudy (anonymous):

Isn't it -5

OpenStudy (anonymous):

you might be right molly?

OpenStudy (anonymous):

so i use the quadratic formula???

OpenStudy (anonymous):

substitute the one by one (-5)^3+7(-5)^2+9(-5)-5 the expression is equal to 0 so x=-5 is a root of the polynomial so now divide the polynomial by (x+5) I think it would be x=-5 -1+sqrt2 -1-sqrt2

OpenStudy (anonymous):

okay......? wait, do I use the quadratic formula for this problem?

OpenStudy (anonymous):

you can use synthetic division (x^3+7x^2+9x-5)/x+5

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i would get x^2+2x-1?

OpenStudy (anonymous):

any expression divided by 1 is the expression so it would be x^2+2x-1*1 then write it as a set of linear factors

OpenStudy (anonymous):

x^2+2x-1*1

OpenStudy (anonymous):

okay. so as a final answer i would get −1±√2

OpenStudy (anonymous):

yes! also -5

OpenStudy (anonymous):

okay. thank you!!!

OpenStudy (anonymous):

x=-5, −1±√2

OpenStudy (anonymous):

thank you so so so much! :)

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

wow good job molly that was impressive!!! @MollyhatesMath @JessilynP This was one of the most helpful threads ive everseen on open study